OpenStudy (anonymous):

Simplify negative 2 and 1 over 6 – negative 7 and 1 over 3.

OpenStudy (anonymous):



OpenStudy (anonymous):

That is how the problem looks

jhonyy9 (jhonyy9):

1 1 - 2-- - 7--- = ? 6 3 is right hence ?

OpenStudy (anonymous):

negative 2 and negative 7

OpenStudy (anonymous):

but yes

jhonyy9 (jhonyy9):

yes there are minus 2 and minus 7

OpenStudy (anonymous):

oh okay then yes

jhonyy9 (jhonyy9):

so there fractions - yes ? how you add or subtract fractions ?

jhonyy9 (jhonyy9):

do you have heared anything about denominator common ?

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

so for 6 and 3 what will be the common denominator ?

jhonyy9 (jhonyy9):

do you know it ?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

wait not 6?

OpenStudy (anonymous):

no*

OpenStudy (anonymous):

@jhonyy9

jhonyy9 (jhonyy9):

yes will be 6 so than what you need to do now ?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

@jhonyy9

jhonyy9 (jhonyy9):

sorry now i have saw it above that you have wrote there your exercise and there is - -7 so this mean that will be minus in front with minus what sign result ?

OpenStudy (anonymous):

im not sure what you mean?

jhonyy9 (jhonyy9):

1 1 - 2 --- - ( - 7 ---- ) = ? 6 3 so this is right now sure

jhonyy9 (jhonyy9):

and this minus (minus 7) what sign what result for in front of 7 ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

minus time minus = ? what sign ?

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

wait no positive

jhonyy9 (jhonyy9):

than now what + or - ?

OpenStudy (anonymous):

-

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

@jhonyy9 where are you getting at with this, i am a bit confused.

jhonyy9 (jhonyy9):

check your exercise how you have wrote above firstly - so there is - -7 --- yes ?

OpenStudy (anonymous):

is the answer -5 1/6 because idk what your talking about?

jhonyy9 (jhonyy9):

but do you know how you get this result ?

OpenStudy (anonymous):

no not really. Can you explain to how?

jhonyy9 (jhonyy9):

1 -2--- = (-(2*6 +1))/6 = - 13/6 6 1 - 7---- = ( - (21+1)/3)) = - 22/3 3 1 - 13/6 - - 22/3 = (- 13 - (-22*2) ) /6= (-13 +44)/6 = 31/6 = 5--- 6 1 so but than you need getting - 5 --- 6 so than this mean that there are in front of 7 just one minus not how you have wrote there doubled the minus sign ok. ?

jhonyy9 (jhonyy9):

@KAKES1967 where are you ?

OpenStudy (anonymous):

the answer was wrong but its fine..

jhonyy9 (jhonyy9):

what answer was wrong ?

OpenStudy (anonymous):

Thanks anyway, the answer -5 1/6 is wrong..

jhonyy9 (jhonyy9):

read what i have wrote to you above please

jhonyy9 (jhonyy9):

so this mean that your exercise you not have wrote it right sure - ok ?

jhonyy9 (jhonyy9):

so than the second term will be just with one minus sign 7 1/3 so than will result your choice result what you have wrote there like right answer - ok ?

jhonyy9 (jhonyy9):

@shaleiah do you understand me please ?

OpenStudy (shaleiah):

Yes, I understand.

OpenStudy (anonymous):

Alright i am done with the question.

jhonyy9 (jhonyy9):

ok. so was my pleasure good luck bye